# PAVLM: Advancing Point Cloud based Affordance Understanding Via Vision-Language Model

### v1

@@ Abstract
Affordance understanding, the task of identifying actionable regions on 3D objects, plays a vital role in allowing robotic systems to engage with and operate within the physical world. Although Visual Language Models (VLMs) have excelled in high-level reasoning and long-horizon planning for robotic manipulation, they still fall short in grasping the nuanced physical properties required for effective human-robot interaction. In this paper, we introduce PAVLM (Point cloud Affordance Vision-Language Model), an innovative framework that utilizes the extensive multimodal knowledge embedded in pre-trained language models to enhance 3D affordance understanding of point cloud. PAVLM integrates a geometric-guided propagation module with hidden embeddings from large language models (LLMs) to enrich visual semantics. On the language side, we prompt Llama-3.1 models to generate refined context-aware text, augmenting the instructional input with deeper semantic cues. Experimental results on the 3D-AffordanceNet benchmark demonstrate that PAVLM outperforms baseline methods for both full and partial point clouds, particularly excelling in its generalization to novel open-world affordance tasks of 3D objects. For more information, visit our project site: pavlm-source.github.io.
@@ I INTRODUCTION
To overcome these limitations, some researchers have proposed the use of VLMs to process point cloud data by projecting it onto depth maps, as demonstrated by PointCLIP [15] and PointCLIP V2 [16], shown in Figure 1. However, these methods involve multiple projection steps, such as quantization, densification, smoothing, and squeezing, which inevitably lead to a loss of geometric and spatial information. In this paper, we introduce a novel framework that directly extracts semantic features from raw point-cloud data. By moving from 2D representations to 3D point clouds, our model preserves the richness of geometric details and ensures a more comprehensive and accurate affordance understanding. Specifically, our proposed point cloud affordance Vision-Language Model (PAVLM) integrates a geometric-guided propagation module, with the semantic reasoning and world knowledge embedded in large language models (LLMs), such as the Llama series [17, 18]. We also prompt Llama-3.1 models to generate refined context-aware text, augmenting instructional input with deeper semantic cues. Experimental results on the 3D-AffordanceNet benchmark demonstrate that our approach significantly improves performance on open-vocabulary affordance learning tasks and shows strong generalization capabilities, particularly in scenarios involving novel objects and unseen environments. By incorporating world knowledge and 3D geometry, our model generalizes more effectively to novel objects, enabling a more intuitive and effective robotic interaction with the 3D world.
@@
  • We propose a unified architecture that integrates 3D point cloud data with an open-source language model for enhanced affordance understanding.
@@ II RELATED WORK
To address this limitation, Copa [38] introduced the Segment Anything Model (SAM), which allows for more precise selection of grasping points, enhancing the generalization abilities of Voxposer. However, since SAM operates only in 2D, it limits planning for robotic manipulation in 3D space. While 2D affordance maps are commonly used, AffordanceLLM [13] was the first to integrate a 2D affordance map with an LLM to improve the agent’s reasoning abilities. Despite this, the gap between 2D affordance reasoning and real-world 3D manipulation remains a challenge. To bridge this gap, we propose the framework of PAVLM to achieve 3D object affordance generation, potentially enabling more effective robotic manipulation in real-world environments.
@@ III The proposed PAVLM Framework
More detailed, the process begins with the geometric-guided point encoder, which extracts geometric features from the input point cloud. These features are then infused into the point cloud embedding to enrich the representation. Next, we utilize the Llama-3.1 LLM model through prompting to enhance the reasoning capabilities of the system via data augmentation, generating more diverse context-aware question-answer pairs. Afterward, we adopt a pre-trained 3D Image-Bind model to align the point cloud embedding with the textual tokens, ensuring multimodal consistency between the 3D data and the instruction. Following this, the fine-tuned Llama-2 model generates the affordance mask labels, which are integrated with the point cloud embeddings. Finally, these embeddings are fed into a 3D affordance decoder, which produces the final affordance map with high precision. A detailed pipeline of our framework is shown in Figure 2.
@@ III-A Geometric-guided Point Encoder
As illustrated in Figure 3, we propose a geometric-guided point encoder to generate the final point cloud embeddings. This encoder is composed of two key submodules: (a) the geometric information extraction module and (b) the feature propagation module. In the first module, we begin by dividing the input point cloud $P\in\mathbb{R}^{2048\times 3}$ into $k$ point patches. Inspired by Point-Bert [29], we applied convolutional layers to extract features from these patches, resulting in point-wise features $\{f_{i}\}_{i=1}^{k}$. These features are then combined with positional embeddings $\{p_{i}\}_{i=1}^{k}$ and fed into a series of transformer blocks via a cross-attention mechanism. This process yields refined geometric features $\{g_{i}\}_{i=1}^{k}$. We adopt a structure comprising 3-layer transformer blocks, and the extraction of geometric features can be formulated as follows:
@@ III-B Textual Instruction Augmentation
To enhance affordance understanding through the alignment of point cloud data and text, we incorporate 3D-specific descriptions that reflect implicit human-object interaction characteristics as the textual input. Rather than using general action terms like cut, handover, and tool use, commonly seen in previous affordance-based robotic works [24, 41], we adopt a more detailed question-answering approach inspired by AffordanceLLM [13]. In this approach, a seed question is generated, such as ”What part of the <object name> should we interact with to <action name> it?” followed by an answer like ”You can <action name> the area <mask token>.”
Given the powerful descriptive capabilities of large language models (LLMs), we extend this approach by utilizing Llama-3.1 70B [42] to generate richer, 3D-specific augmentation question-answering pairs. Llama-3.1 processes the language command and outputs responses based on its pretrained knowledge, resulting in more detailed and context-aware interactions. An example of the prompt we used for this process is illustrated in Figure 4. For a 3D dataset with $K$ categories, we replace the <mask label> in each command with the corresponding category value and input these commands into Llama-3.1, generating 3D-specific descriptions enriched with category-specific semantics.
@@ III-C Feature Alignment of Point Cloud and Text
where $\textit{eps}\cdot e$ term is used to avoid numerical instability or division by zero in distance calculations.
@@ III-D Masked-based Affordance Decoding
Seen in Figure 3, this query embedding $q$ first concatenates with the original point cloud $P$, and then integrated with point cloud embedding $P_{em}$ into the 3D affordance decoder, which consists of two convolutional layers and a batch normalization layer. Its primary function is to reduce the dimensionality of the multimodal feature embeddings and produce a point-wise affordance map $\mathbf{M}\in\mathbb{R}^{2048\times 1}$ in the point cloud, where each point $M_{i}\in[0,1]$ represents the affordance score for point $i$ in the point cloud. The above process can be mathematically expressed as:
@@ IV-A Experiment Setup
Metrics. We assess affordance estimation using four metrics: mean Average Precision (mAP), Area Under the ROC Curve (AUC), average Intersection Over Union (aIoU), and mean squared error (MSE). Firstly, we generate a Precision-Recall curve and compute the Average Precision for each affordance. The area under the ROC curve is calculated as the AUC metric. We also compute the mean squared error (MSE) for each affordance class and sum the results across all categories. After binarizing the predictions in a step of 0.01, the aIoU is calculated by progressively increasing the threshold from 0 to 0.99. Aside from MSE, all metrics are further process In an average mathematical operation.
Baselines. We compare our approach with state-of-the-art baseline methods. As shown in Figure 1, we mainly compare our work with PointCLIP [15] and PointCLIP V2 [16], which both achieve great generalization performance on the language-interactive 3d point cloud understanding tasks.
  • PointCLIP [15]: PointCLIP is the first model to bridge the gap between the 2D image-text pair pre-trained model (CLIP) and the irregular structure of 3D point clouds. It introduces a multi-view projection-based method that transform point clouds into multi-view depth maps, which are then processed by CLIP’s visual encoder. It achieves promising results on standard 3D benchmarks such as ModelNet and ScanObjectNN.
  • PointCLIP V2 [16]: PointCLIP V2 is an enhanced framework that integrates CLIP and GPT models for powerful 3D open-world learning. Building on the original PointCLIP, PointCLIP V2 introduces more realistic visual projections and rich 3D-specific text prompts, making it a robust and versatile tool for open-world point cloud understanding.
@@
Dataset. We primarily evaluate our methods using the 3D AffordanceNet benchmark [43], which is specifically designed for visual affordance understanding in 3D. The dataset contains 22,949 3D object models across 23 semantic categories and covers 18 affordance categories. During the training process, we employ two dataset-splitting strategies: seen and unseen. For the seen strategy, we shuffle the original dataset and split it into training, validation, and test sets using an 8:1:1 ratio where each of these splits contains all categories. In contrast, the unseen strategy follows the same splitting ratio but introduces specific object categories, namely ”mug,” ”knife,” and ”scissors”, which are entirely reserved for the test set. These categories are excluded from the training and validation sets, making this split particularly valuable for evaluating the model’s ability to generalize to novel objects, which is significant for real-world robotic manipulation tasks.
@@ IV-B Ablation Study
We performed several ablation studies to evaluate how various modules of our model contribute to its overall performance on the seen split dataset. Specifically, we examined the effects of varying text prompts and point cloud encoders. Following the evaluation prompts used in previous work [2], we tested four different types of text prompts to analyze their impact on model performance: (1) A simple greeting noise, such as ”Hi”, (2) The action associated with the object, (3) A combination of the object name and action label, and (4) A full question prompt and its augmentation, such as ”Which specific region of the mug should we target to provide drinking?”.
@@
In addition, we compared our geometric-guided point encoder against two widely used point cloud-based network architectures: PointNet++ [25] and DGCNN [40]. As demonstrated in Table II, our proposed geometric-guided point encoder average outperforms them when used as the vision backbone. For instance, in key metrics like mAP and aIoU, our model achieves over 10% higher performance than the best PointNet++ model. This improved performance indicates that our encoder is more effective in extracting meaningful semantic feature embeddings for our affordance understanding task.
@@ IV-C Evaluation Result
Comparison with Baselines. To further evaluate 3D affordance understanding, we compared the performance of our model with state-of-the-art point cloud understanding approaches, specifically PointCLIP [15] and PointCLIP V2 [16]. Since both baseline methods leverage the frozen pre-trained vision-language model CLIP, we introduced a variant of our model by freezing the pre-trained Image-Bind and Llama-2 modules. The results presented in Table III, indicate that both baseline methods struggle with our 3D affordance learning task. In contrast, our model, including the frozen variant, consistently delivers superior performance across all four metrics, demonstrating its effectiveness on both full-shape and partial-shape point cloud data.
@@ IV-D Visualization Result
As illustrated in Fig 5, we present several qualitative results from our model on unseen objects and categories. When given an implicit instruction, such as ”Where on the chair should we use it for sitting?”, our model accurately identifies the corresponding affordance region on the chair. In some categories from the unseen split, like knives, more complex instructions that require extensive reasoning ability, such as ”To use the knife as a tool, which area should we interact with?”, are handled effectively by our model. It accurately predicts the handle centre of the knife. These results demonstrate that our method leverages the reasoning capabilities of LLMs to generate high-quality affordance masks, even for challenged partial point cloud understanding.
V CONCLUSIONS

### v2

@@ Abstract
Affordance understanding, the task of identifying actionable regions on 3D objects, plays a vital role in allowing robotic systems to engage with and operate within the physical world. Although Visual Language Models (VLMs) have excelled in high-level reasoning and long-horizon planning for robotic manipulation, they still fall short in grasping the nuanced physical properties required for effective human-robot interaction. In this paper, we introduce PAVLM (Point cloud Affordance Vision-Language Model), an innovative framework that utilizes the extensive multimodal knowledge embedded in pre-trained language models to enhance 3D affordance understanding of point cloud. PAVLM is an approach to integrates a geometric-guided propagation module with hidden embeddings from large language models (LLMs) to enrich visual semantics. On the language side, we prompt Llama-3.1 models to generate refined context-aware text, augmenting the instructional input with deeper semantic cues. Experimental results on the 3D-AffordanceNet benchmark demonstrate that PAVLM outperforms baseline methods for both full and partial point clouds, particularly excelling in its generalization to novel open-world affordance tasks of 3D objects. For more information, visit our project site: pavlm-source.github.io.
@@ I INTRODUCTION
To overcome these limitations, some researchers have proposed the use of VLMs to process point cloud data by projecting it onto depth maps, as demonstrated by PointCLIP [15] and PointCLIP V2 [16], shown in Figure 1. However, these methods involve multiple projection steps, such as quantization, densification, smoothing, and squeezing, which inevitably lead to a loss of geometric and spatial information. ULIP [17] and ULIP2 [18] have tried to align the point cloud with the image and text that is integrated in LLM as 3D-LLM [19] and which will be able to perform the point cloud understanding or point cloud generation task. In this paper, we introduce a framework that integrates knowledge from large language models (LLMs) with 3D point cloud data to enhance point-to-point affordance understanding. We prompt Llama-3.1 models to generate refined, context-aware text, thereby augmenting instructional input with deeper semantic cues. Experimental results on the customized 3D-AffordanceNet benchmark demonstrate that our approach significantly improves performance on open-vocabulary affordance learning tasks and exhibits strong generalization capabilities, particularly in scenarios involving novel objects and unseen environments.
@@
  • We propose a unified architecture that integrates 3D point cloud data with an open-source language model for point cloud affordance understanding, especially considering the real-robot partial views.
@@ II RELATED WORK
To address this limitation, Copa [39] introduced the Segment Anything Model (SAM), which allows for more precise selection of grasping points, enhancing the generalization abilities of Voxposer. However, since SAM operates only in 2D, it limits planning for robotic manipulation in 3D space. While 2D affordance maps are commonly used, AffordanceLLM [13] was the first to integrate a 2D affordance map with an LLM to improve the agent’s reasoning abilities. Despite this, the gap between 2D affordance reasoning and real-world 3D manipulation remains a challenge. To address this, approaches such as LASO [40] have leveraged language output as cues for 3D point cloud feature reconstruction. However, these methods do not fully exploit the multimodal capabilities of pre-trained large language models that are already aligned with point cloud data. To overcome this limitation, we propose PAVLM, a framework designed to generate 3D object affordances, potentially enabling more effective robotic manipulation in real-world environments.
@@ III The proposed PAVLM Framework
More detailed, the process begins with the geometric-guided point encoder, which extracts geometric features from the input point cloud. These features are then infused into the point cloud embedding to enrich the representation. Next, we utilize the Llama-3.1 LLM model through prompting to enhance the reasoning capabilities of the system via data augmentation, generating more diverse context-aware question-answer pairs. Afterward, we adopt a pre-trained 3D Image-Bind model to align the point cloud embedding with the textual tokens, ensuring multimodal consistency between the 3D data and the instruction. Subsequently, the fine-tuned Llama-2 model generates affordance mask labels, which we integrate with the point cloud embeddings. We selected Llama-2 due to the limitations of LLaMA-Adapter v1 [41], while still benefiting from its ability to incorporate point-cloud knowledge, as seen in ImageBind-LLM[33]. Furthermore, pre-trained Llama-2 knowledge supports our affordance task sufficiently without requiring extensive language-specific fine-tuning. Thus, we focus primarily on fine-tuning it for affordance map label classification. Finally, these embeddings are fed into a 3D affordance decoder, which produces the final affordance map with high precision. A detailed pipeline of our framework is shown in Figure 2.
@@ III-A Geometric-guided Point Encoder
As illustrated in Figure 3, we propose a geometric-guided point encoder to generate the final point cloud embeddings. This encoder is composed of two key submodules: (a) the geometric information extraction module and (b) the feature propagation module. In the first module, we begin by dividing the input point cloud $P\in\mathbb{R}^{2048\times 3}$ into $k$ point patches. Inspired by Point-BERT [30], we define a constant K to form K groups, each containing 256 points. We then apply convolutional layers to these patches to extract point-wise features $\{f_{i}\}_{i=1}^{k}$. The features are added to the positional embeddings $\{p_{i}\}_{i=1}^{k}$ and fed into a series of transformer blocks via a cross-attention mechanism. This process yields refined geometric features $\{g_{i}\}_{i=1}^{k}$. We adopt a structure comprising 3-layer transformer blocks, and the extraction of geometric features can be formulated as follows:
@@ III-B Textual Instruction Augmentation
To enhance affordance understanding through the alignment of point cloud data and text, we incorporate 3D-specific descriptions that reflect implicit human-object interaction characteristics as the textual input. Rather than using general action terms like cut, handover, and tool use, commonly seen in previous affordance-based robotic works [25, 44], we adopt a more detailed question-answering approach inspired by AffordanceLLM [13]. In this approach, a seed question is generated, such as “What part of the <object name> should we interact with to <action name> it?” followed by an answer like “You can <action name> the area <mask token>”.
Given the powerful descriptive capabilities of large language models (LLMs), we extend this approach by using Llama-3.1 70B [45] to generate richer, 3D-specific augmentation question-answering pairs. Llama-3.1 processes the language command and outputs responses based on its pre-trained knowledge, resulting in more detailed and context-aware interactions. An example of the prompt we used for this process is illustrated in Figure 4. For a 3D dataset with $K$ categories, we replace the <mask label> in each command with the corresponding category value and input these commands into Llama-3.1, generating 3D-specific descriptions enriched with category-specific semantics.
@@ III-C Feature Alignment of Point Cloud and Text
where $\textit{eps}\cdot e$ the term is used to avoid numerical instability or division by zero in distance calculations.
@@ III-D Masked-based Affordance Decoding
Seen in Figure 3, this query embedding $q$ first concatenates with the original point cloud $P$, and then integrated with point cloud embedding $P_{em}$ into the 3D affordance decoder, which consists of two convolutional layers and a batch normalization layer. Its primary function is to reduce the dimensionality of the multimodal feature embeddings and produce a point-wise affordance map $\mathbf{M}\in\mathbb{R}^{2048\times 1}$ in the point cloud, where each point $M_{i}\in[0,1]$ represents the affordance score for a point $i$ in the point cloud. The above process can be mathematically expressed as:
@@ IV-A Experiment Setup
Metrics. We assess affordance estimation using four metrics: mean Average Precision (mAP), Area Under the ROC Curve (AUC), average Intersection Over Union (aIoU), and mean squared error (MSE). Firstly, we generate a Precision-Recall curve and compute the Average Precision for each affordance. The area under the ROC curve is calculated as the AUC metric. We also compute the mean squared error (MSE) for each affordance class and sum the results across all categories. After binarizing the predictions in a step of 0.01, the aIoU is calculated by progressively increasing the threshold from 0 to 0.99. Aside from MSE, all metrics are further processes in an average mathematical operation. All numerical values reported in the tables have been scaled by a factor of 100 (i.e., multiplied by 100) and values are rounded at the last reported digit.
Baselines. We compare our approach with state-of-the-art baseline methods. As shown in Figure 1, we mainly compare our work with PointCLIP [15] PointCLIP V2 [16], ULIP [17], and ULIP2 [18] which all achieve great generalization performance on the language-interactive 3d point cloud understanding tasks.
  • PointCLIP and PointCLIP V2 [15, 16]: PointCLIP bridges the gap between 2D image-text pre-trained models (CLIP) and the irregular structure of 3D point clouds by introducing a multi-view projection-based method. This approach transforms point clouds into multi-view depth maps, which are then processed by CLIP’s visual encoder, achieving promising results on standard 3D benchmarks such as ModelNet and ScanObjectNN. Building upon this, PointCLIP V2 enhances the framework by integrating CLIP and GPT models for robust 3D open-world learning. It introduces more realistic visual projections and rich 3D-specific text prompts, making it a versatile tool for open-world point cloud understanding.
  • ULIP and ULIP-2 [17, 18]: ULIP is a multimodal pre-training framework that aligns 3D point clouds, images, and textual descriptions into a unified representation space. By leveraging pre-trained vision-language models, ULIP enhances 3D representation learning, enabling state-of-the-art performance in tasks like zero-shot 3D classification and image-to-3D retrieval. Extending this approach, ULIP-2 employs large multimodal models to automatically generate comprehensive language descriptions for 3D objects, facilitating scalable multimodal pre-training without manual annotations. This advancement further improves 3D understanding and cross-modal applications.
@@
Dataset. We primarily evaluate our methods using the 3D AffordanceNet benchmark [46], which is specifically designed for visual affordance understanding in 3D. The dataset contains 22,949 3D object models across 23 semantic categories and covers 18 affordance categories. During the training process, we employ two dataset-splitting strategies: seen and unseen. For the seen strategy, we shuffle the original dataset and split it into training, validation, and test sets using an 8:1:1 ratio where each of these splits contains all categories. In contrast, the unseen strategy follows the same splitting ratio but introduces specific object categories, namely “mug,” “knife,” and “scissors”, which are entirely reserved for the test set. These categories are excluded from the training and validation sets, making this split particularly valuable for evaluating the model’s ability to generalize to novel objects, which is significant for real-world robotic manipulation tasks. We have further processed the data into rows containing the fields “instruct”, “input”, “Answer”, and “Affordance_map” to facilitate downstream tasks. The original template for the “instruct” field, inspired by AffordanceLLM [47], is formulated as: “What part of the {semantic_class} should we interact with in order to {affordance} it?” Correspondingly, the “Answer” field is structured as: “You can {affordance} the area {mask_token}.” The “instruct” and “answer” field is further augmented using the method described in Section III-B.
@@ IV-B Ablation Study
We performed several ablation studies to evaluate how various modules of our model contribute to its overall performance on the seen split dataset. Specifically, we examined the effects of varying text prompts and point cloud encoders. Following the evaluation prompts used in previous work [2], we tested four different types of text prompts to analyze their impact on model performance: (1) A simple greeting noise, such as “Hi”, (2) The action associated with the object, (3) A combination of the object name and action label, and (4) A full question prompt and its augmentation, such as “Which specific region of the mug should we target to provide drinking?”.
@@
In addition, we compared our geometric-guided point encoder against two widely used point cloud-based network architectures: PointNet++ [26] and DGCNN [43]. As demonstrated in Table II, our proposed geometric-guided point encoder average outperforms them when used as the vision backbone. For instance, in key metrics like mAP and aIoU, our model achieves over 10% higher performance than the best PointNet++ model. The main reason is that feature propagation provides sufficient cues to the upsampling mechanism, especially in partial shapes, so that before classification, there is already a strong understanding of the object. The framework allows the mask label to trigger the activation part more effectively, resulting in better performance. This enhanced performance indicates that our encoder is more effective in extracting meaningful semantic feature embeddings for our affordance understanding task.
@@ IV-C Evaluation Result
Comparison with Baselines. To further evaluate 3D affordance understanding, we compared the performance of our model with state-of-the-art point cloud understanding approaches, specifically PointCLIP [15], PointCLIP V2 [16], ULIP [17], and ULIP2 [18]. Since all baseline methods leverage the frozen pre-trained vision-language model such as PoinCLIP from CILP and ULIP have reported using Pointbert as their core point encoder would perform the best, we introduced a variant of our model by freezing the pre-trained Image-Bind and Llama-2 modules, then further fine-tune by our dataset. It is noted that previous baseline methods typically achieve point cloud understanding using full-shape point clouds from a computer vision perspective. In contrast, we emphasize real-robot visual understanding by introducing extra partial-shape point clouds. The results presented in Table III are obtained using the original methods and their best-released checkpoints, fine-tuned for three epochs, and essential data transformations to fit the input size and model requirements. Note that the PointCLIP and ULIP series were originally trained on other tasks in their original repository, which we assume gave them general capabilities for understanding the point cloud in downstream tasks. In contrast, our model, including the frozen variant, consistently delivers superior performance across all four metrics, demonstrating its effectiveness on both full-shape and partial-shape point cloud data.
@@
Full-shape ULIP 7.6 50.0 0.4 0.60
Full-shape ULIP2 7.6 50.1 1.2 0.63
@@
Partial-shape ULIP 9.2 50.0 1.0 0.58
Partial-shape ULIP2 9.2 50.1 1.1 0.55
@@ IV-D Visualization Result
As illustrated in Fig 5, we present several qualitative results from our model on unseen objects and categories. When given an implicit instruction, such as “Where on the chair should we use it for sitting?”, our model accurately identifies the corresponding affordance region on the chair. In some categories from the unseen split, like knives, more complex instructions that require extensive reasoning ability, such as “To use the knife as a tool, which area should we interact with?”, are handled effectively by our model. It accurately predicts the handle centre of the knife. These results demonstrate that our method leverages the reasoning capabilities of LLMs to generate high-quality affordance masks, even for challenged partial point cloud understanding.
V Quantitative Limitations
While our proposed framework, PAVLM, demonstrates significant advancements in 3D affordance understanding by integrating geometric-guided propagation with large language models (LLMs), certain quantitative limitations must be acknowledged. Due to time constraints and overlapping research in the visual domain, we were unable to perform direct, method-wise comparisons with similar work, such as LASO [40] and 3D-LLM [19]. However, a theoretical comparison can be explained. 3D-LLM employs an approach similar to that of PointCLIP by extracting multi-view images from the 3D scene, deriving dense 2D features, and concatenating them to reconstruct 3D representations—a process that inevitably leads to some loss of the original 3D information. In contrast, LASO does not leverage 3D embeddings to fully exploit the multimodal capabilities of pre-trained LLMs, which may limit its performance in open-world recognition tasks involving unseen categories compared to PAVLM.
VI CONCLUSIONS